# Reasoning Efficiently Through Adaptive Chain-of-Thought Compression: A Self-Optimizing Framework

### v1

@@ 3.1.3. Evaluation Metrics
Accuracy. We use Pass@1 as the primary metric to evaluate correctness. This metric calculates the proportion of problems for which the model produces a correct solution on the first attempt, directly reflecting its practical effectiveness in solving tasks.
[…]
Average Token Length. We report the average number of tokens generated in model outputs. This metric reflects the verbosity of the reasoning process and helps assess whether excessively long outputs introduce redundancy or cause truncation. It also facilitates a direct comparison of output lengths between models with and without CoT reasoning.
3.1.4. Implementation Details
[…]
3.2. Empirical Findings
[…]
Truncation Leads to Performance Degradation. Table 1 demonstrates that excessively long CoT outputs introduce severe truncation issues, which directly limit performance improvements. For the smallest model (DeepSeek-R1-Distill-Qwen-1.5B), accuracy drops sharply, with the pass@1 score falling from 60.36% to 46.34%. This loss coincides with an extreme average CoT length of 7,645 tokens. The underlying cause is frequent truncation, which removes critical parts of the solution. In fact, 63 out of 164 problems are truncated on this model. As the model size increases, the truncation problem becomes less frequent, and moderate gains emerge. The 7B and 14B models show performance gains with CoT, but these improvements are limited by long outputs and truncation. The 32B model still experiences truncation in about 8.5% of cases, which keeps its pass@1 unchanged at 90.24% with or without CoT.
Table 1. HumanEval Results of Reasoning Models (With vs. Without CoT)
Models HumanEval (With CoT) HumanEval (Without CoT)
Pass@1 (%) Avg Tokens Trunc Loop Pass@1 (%) Avg Tokens Trunc Loop
DeepSeek-R1-Distill-Qwen-1.5B 46.34 7,645 63 62 60.36 513 0 0
DeepSeek-R1-Distill-Qwen-7B 80.49 4,210 24 22 70.12 474 0 0
DeepSeek-R1-Distill-Qwen-14B 89.63 3,687 16 14 84.76 444 0 0
DeepSeek-R1-Distill-Qwen-32B 90.24 3,340 14 12 90.24 474 0 0
[…]
Finding 1:  Excessive CoT length induces harmful truncation, especially in small-scale LLMs. Small-scale LLMs are more easily affected by long CoTs because their context is quickly filled. For larger models, truncation is less frequent but still reduces the benefit, showing that longer reasoning does not always improve results once the context limit is reached.
[…]
Reasoning loop results in truncation. We manually examine truncation cases to better understand their impact on model performance. Our analysis shows that the model often falls into repetitive reasoning loops, repeatedly generating the same fragment. For example, it may continually output phrases such as “Wait, if …” until the maximum context length is reached. This behavior consumes a large number of tokens and severely reduces efficiency. As shown in Table 1, reasoning loops account for most truncation cases, with an average rate of 94.0%. When CoT generation is disabled, this problem disappears entirely. Larger models alleviate the issue to some extent, but reasoning loops are not fully eliminated.
These observations suggest that while CoT can enhance reasoning, uncontrolled generation often leads to loops and truncation, which harm both efficiency and reliability. Therefore, controlling CoT length is crucial to preserve its benefits without incurring instability.
[…]
Finding 2:  Reasoning loops are the dominant reason for truncation. They waste tokens, reduce efficiency, and persist across model scales, disappearing only when CoT is disabled. This highlights looping as a critical issue in CoT reasoning.
[…]
Is longer always better? While prior work (Jin et al., 2024) suggests that increasing the length of CoT reasoning improves accuracy, our findings reveal a different pattern. We conduct experiments on one of the most challenging problems in HumanEval (findmyway, 2024), namely HumanEval/129. Since the 1.5B model achieves almost no correct solutions in our experiments (pass@1 is nearly zero), we exclude it from further analysis. For each of the other models, we generate 10,000 responses and examine the distribution of CoT lengths, as shown in Figure 1, with the maximum token length set to 16,384. Contrary to the common assumption that longer reasoning consistently leads to better performance (Jin et al., 2024), our analysis reveals a clear pattern across all three models. Failed outputs are frequently associated with longer CoT lengths, while successful outputs are more concentrated around shorter lengths. For instance, in the 7B model, the median CoT length of failed generations (9,489 tokens) surpasses that of successful generations (8,296 tokens), a pattern that persists in both the 14B and 32B models. This suggests that excessively long reasoning chains may introduce redundancy, amplify error propagation, or even trigger context truncation, ultimately degrading performance. These results highlight that effective reasoning does not simply require “more thinking”, but rather appropriate reasoning lengths that balance sufficiency and conciseness.
[…]
Finding 3:  Longer CoT does not guarantee higher accuracy. Across all three models, failed generations consistently exhibit longer CoT lengths than successful ones (Figure 1). This indicates that excessively long reasoning chains often introduce redundancy and error propagation rather than improving performance, highlighting the need for appropriately controlled CoT lengths.
[…]
Long CoT introduces application latency. Excessively long CoT reasoning introduces significant latency overhead, which ultimately degrades user experience. To quantify this impact, we evaluate the inference performance of Deepseek-R1-Distill-Qwen models of varying sizes using the transformers library on a single NVIDIA A800 GPU. The measured generation speeds range from 46 tokens/s for the smallest (1.5B) model to 18 tokens/s for the largest (32B) model. Using these measurements, we estimate the average end-to-end latency on two benchmarks: HumanEval, representing relatively simple coding tasks, and Codeforces, which contains more complex algorithmic challenges. The results are presented in Figure 2.
Figure 2. Inference latency of different model sizes on HumanEval and Codeforces datasets with and without CoT reasoning. While CoT improves reasoning capability, it substantially increases application latency, particularly for larger models, posing practical constraints on real-world deployment.
Figure 2 shows that incorporating CoT substantially increases response time, particularly for larger models on complex tasks. On Codeforces, the 32B model with CoT requires an average of 776 seconds to complete the reasoning for each problem, compared to 157 seconds without CoT. Even the smaller 1.5B model takes more than five minutes per problem when using CoT. These delays make real-time use impractical. Moreover, all experiments were conducted under a 16k-token context limit, whereas Codeforces samples already exceed 14k tokens on average, implying that real-world latency would be even higher. Thus, while CoT improves reasoning, its verbosity imposes severe constraints on latency-sensitive applications, highlighting the need for effective CoT length control.
[…]
Finding 4:  Long CoT substantially slows down inference. For example, the 32B model on Codeforces requires nearly five times longer with CoT than without. This latency overhead makes CoT impractical for real-time or latency-sensitive applications, underscoring the need for effective length control.
@@ 4.1. Pre-Inference Data Generation
The first stage of SEER focuses on constructing a comprehensive corpus of CoT outputs to serve as the foundation for subsequent filtering and fine-tuning. In this step, the base language model is used to process each question in the training set, generating a complete solution that includes an explicit, step-by-step reasoning path.
A distinctive advantage of our approach is its low sensitivity to prompt engineering. While a prompt is required to elicit CoT reasoning, our subsequent adaptive filtering stage ensures that the method is not critically dependent on the choice of prompt format. This property allows us to adopt a pragmatic strategy: whenever an official, dataset-specific prompt template is available, we use it directly; otherwise, we revert to a robust general-purpose template. For example, for mathematical reasoning tasks, we utilize the system prompt from the study by Wei et al. (Wei et al., 2022):
@@
To prevent incomplete reasoning caused by premature output truncation, we allocate a sufficiently high maximum token budget during generation. This ensures that the model’s natural CoT length distribution is faithfully captured. Accurately profiling this distribution is critical for calibrating the adaptive filter in later stages, as it allows us to set dataset-specific compression thresholds that minimize information loss.
4.2. Best-of-N Sampling for Data Refinement
To prevent the reasoning loops noted in Section 3, which can harm data quality, we adopt a BoN sampling strategy to build a high-quality training set. This strategy also helps obtain cleaner datasets with more concise CoT lengths. For each input question, the model generates $\mathcal{N}$ candidate responses, each augmented with a CoT. These candidates are then ranked and filtered according to a strict hierarchical set of criteria:
  1. (1)
    Correctness of the Final Answer: Only candidates producing the correct final answer are considered. All incorrect outputs are discarded.
  2. (2)
    Presence of Valid CoT: From the remaining candidates, we retain only those containing a non-empty reasoning path. This guarantees that each example is explainable and preserves its interpretability.
  3. (3)
    Conciseness of Reasoning: If multiple candidates satisfy the first two criteria, we select the one with the shortest CoT length. This prioritizes efficiency and prevents the inclusion of unnecessarily verbose reasoning.
This multi-stage filtering process ensures that the resulting dataset consists exclusively of reasoning traces that are (1) correct in outcome, (2) logically explicit, and (3) concise. By integrating BoN sampling into our pipeline, we are able to suppress reasoning loops, improve data quality, and provide the model with more effective fine-tuning signals.
4.3. Adaptive CoT Filtering
Even after BoN sampling, CoT responses can vary widely in length depending on the task, domain, and problem complexity. Excessively long CoTs waste computational resources, while overly aggressive compression risks discarding essential reasoning steps. To balance these factors, we introduce an adaptive CoT filtering mechanism that sets a dataset-specific maximum CoT length, denoted $\lambda_{c}$.
A fixed, manually chosen $\lambda_{c}$ is often suboptimal because optimal CoT lengths differ substantially across domains (e.g., mathematical proofs vs. code debugging). If $\lambda_{c}$ is too low, important intermediate reasoning may be lost, potentially reducing accuracy and even causing the model to overfit to overly compressed reasoning patterns. Conversely, a value set too high will dilute the benefits of compression, leading to minimal gains in efficiency.
Our adaptive filtering addresses this by first generating CoT responses for the entire training set during the pre-inference stage and then computing two statistical measures of their length distribution: the mean length $\bar{\lambda}$ and the median length $\tilde{\lambda}$. We define $\lambda_{ada}$ as:
$\lambda_{ada}=\frac{\bar{\lambda}+\tilde{\lambda}}{2}$
This formula anchors the threshold to the observed characteristics of each dataset, balancing robustness (median) with representativeness (mean).
While $\lambda_{c}$ provides a principled baseline, our design remains flexible. Users can manually adjust it to match specific efficiency or accuracy requirements. This tunable control allows practitioners to trade off between computational savings and reasoning completeness, enabling SEER to adapt seamlessly across diverse reasoning scenarios.
4.4. Overall Procedure
Input: Training set $\mathcal{D}=\{(x_{i},y_{i}^{\star})\}$, base model $\mathcal{M}$, candidates per input $\mathcal{N}$
Output: The model $\mathcal{M}_{\text{adapt}}$ fine-tuned on the filtered dataset $\widetilde{\mathcal{D}}$
Stage 1: Pre-inference generation;
foreach $(x_{i},y_{i}^{\star})\in\mathcal{D}$ do
  Generate $\mathcal{N}$ CoT-augmented candidates with $\mathcal{M}$ under a fixed prompt;
  $r_{ij}\leftarrow\text{final response generated for }x_{i}$;
  $\text{CoT}_{ij}\leftarrow\text{reasoning trace (Chain-of-Thought) leading to }r_{ij}$;
  $\mathcal{C}_{i}\leftarrow\{\,(r_{i1},\text{CoT}_{i1}),\ldots,(r_{i\mathcal{N}},\text{CoT}_{i\mathcal{N}})\,\}$;
Stage 2: Best-of-N (BoN) sampling;
foreach $(x_{i},y_{i}^{\star}),\mathcal{C}_{i}$ do
  $\mathcal{S}_{i}\leftarrow\{(r,\text{CoT})\in\mathcal{C}_{i}\mid\textsc{Correct}(r,y_{i}^{\star})\ \wedge\ |\text{CoT}|>0\}$ if $\mathcal{S}_{i}\neq\emptyset$ then
     $(r_{i}^{\star},\text{CoT}_{i}^{\star})\leftarrow\arg\min_{(r,\text{CoT})\in\mathcal{S}_{i}}|\text{CoT}|$ Add $(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})$ to $\mathcal{D}^{\prime}$;
Stage 3: Adaptive CoT filtering;
Compute length set $L\leftarrow\{\,|\text{CoT}_{i}^{\star}|\mid(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})\in\mathcal{D}^{\prime}\,\}$;
$\bar{\lambda}\leftarrow\text{mean}(L),\ \tilde{\lambda}\leftarrow\text{median}(L)$;  $\lambda_{c}\leftarrow(\bar{\lambda}+\tilde{\lambda})/2$;
$\widetilde{\mathcal{D}}\leftarrow\{(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})\in\mathcal{D}^{\prime}\mid|\text{CoT}_{i}^{\star}|\leq\lambda_{c}\}$;
Stage 4: Fine-tuning;
Fine-tune model on $\widetilde{\mathcal{D}}$ with SFT to obtain $\mathcal{M}_{\text{adapt}}$;
return $\mathcal{M}_{\text{adapt}}$;
Algorithm 1 Self-Enhancing Training Process with BoN & Adaptive CoT Filtering
Algorithm 1 summarizes the end-to-end process, from pre-inference generation to BoN-based selection, adaptive filtering, and final fine-tuning on the curated dataset.
Through the combination of high-coverage pre-inference generation, strict BoN filtering, and adaptive CoT length control, SEER produces training data that is both high-quality and efficiency-oriented. This not only improves the reasoning efficiency of the fine-tuned model but also mitigates failure modes such as infinite reasoning loops and truncation.
@@ 5.1. Datasets
To thoroughly evaluate the performance and efficacy of our method, we conduct experiments across three distinct software engineering tasks and one mathematical reasoning task. These tasks were selected to cover a broad spectrum of problem types, input–output modalities, and reasoning requirements. By incorporating both code-related and mathematical reasoning benchmarks, we aim to investigate not only the in-domain effectiveness of our method but also its adaptability to different reasoning paradigms and task formats.
  • MathQA-Python (Austin et al., 2021) is a dataset designed for generating Python programs that solve mathematically formulated problems. Each problem is presented in natural language and requires the model to produce an executable Python solution. The challenges in this dataset include parsing problem statements, performing intermediate symbolic reasoning, and implementing the reasoning steps as correct Python code. This dataset serves as a baseline for evaluating code generation in mathematically oriented problem-solving contexts.
  • CodeXGLUE-Defect-Detection (Zhou et al., 2019) is a binary classification dataset that assesses a model’s ability to identify whether a given piece of source code contains security vulnerabilities or functional defects. The data originates from real-world code bases, and the detection task often requires understanding complex control flows, data dependencies, and security implications. This dataset evaluates the model’s capability in static code analysis and secure software engineering scenarios, where concise yet accurate reasoning is critical to avoid excessive inference overhead.
  • Code-Search-WebQuery (Lu et al., 2021; Husain et al., 2019) is used to test semantic code retrieval capabilities. In this task, the model is provided with a natural language query and a code snippet, and must determine whether the code correctly satisfies the described requirement. This problem is framed as binary classification, and it challenges the model to bridge the semantic gap between natural language descriptions and programming logic. The task reflects realistic software development workflows, such as searching for relevant functions in large code bases.
  • GSM-8K (Cobbe et al., 2021) is a large-scale benchmark for multi-step mathematical reasoning. Each question requires systematic decomposition into intermediate steps before arriving at the final answer. Unlike MathQA-Python, the output here is not executable code but a final numeric answer, making it an ideal testbed for evaluating whether our method can compress reasoning chains without degrading accuracy in purely symbolic reasoning scenarios. It also provides an out-of-domain evaluation for models fine-tuned on code-related tasks.
Additionally, in RQ3 we include the HumanEval dataset introduced in Section 3 to evaluate the generalizability of our method. By selecting this combination of datasets, our evaluation covers tasks ranging from code synthesis, defect detection, and semantic code retrieval to mathematical reasoning.
@@ 5.2. Implementation Details
All experiments are conducted on a server equipped with four NVIDIA RTX A800 GPUs (80 GB each) and 128 Intel Xeon CPUs, running Ubuntu 22.04 LTS. We fine-tune DeepSeek-R1-Distill-Qwen-7B model (DeepSeek-AI, 2025b) under a full-parameter supervised fine-tuning (SFT) setting, using the LLaMAFactory framework (Zheng et al., 2024). Training is performed for three epochs with an effective batch size of eight, gradient checkpointing enabled, and the AdamW optimizer. The learning rate is fixed at $1\times 10^{-5}$ with a cosine decay schedule and 10% warm-up, and gradient clipping (norm 1.0) is applied for stability. During fine-tuning, inputs are tokenized with a maximum sequence length of 8,192 tokens, truncating from the input side if necessary to preserve the output reasoning chain. Although we adopt full-parameter SFT by default, the SEER framework is also compatible with parameter-efficient fine-tuning methods such as LoRA and QLoRA, which we discuss in the Discussion section.
@@ 5.3. Baselines
  • Naive BoN (Munkhbat et al., 2025b): Use only the Best-of-N method to process the pre-inference data and then train the model.
In order to ensure the fairness of the experiment, we have implemented several modifications to the TokenSkip source code. These changes include: 1) the incorporation of the “deepseek” prompt template, as this template is employed by the chosen model; and 2) the use of the vllm library in place of the transformers library to achieve more efficient inference and to mitigate any potential precision loss that could lead to unfair comparisons.
@@ 5.4. Evaluation Metrics
[…]
Accuracy: We use the same pass@1, which is mentioned in Section 3 as the primary metric to evaluate correctness. For programming tasks (e.g., MathQA-Python), a prediction is considered correct if the generated code passes all test cases. For mathematical reasoning tasks (e.g., GSM-8K), we adopt exact matching criterion, where the model’s final answer must exactly match the ground truth.
[…]
Compression Rate: We define the Compression Rate $\rho$ to quantify the conciseness of the generated CoT. Specifically, $\rho$ measures the relative reduction in the number of tokens compared to a reference CoT. It is computed as: $\rho=(\lambda-{\lambda^{\prime}})/{\lambda}$
In this formula, $\lambda$ denotes the token count of the reference (original) CoT, and $\lambda^{\prime}$ represents the token count of the CoT generated by the evaluated model or baseline. A higher value of $\rho$ indicates a greater degree of compression, corresponding to a more concise reasoning path. This metric is particularly important for evaluating methods such as TokenSkip and Naive BoN, which are explicitly designed to produce shorter CoTs without compromising correctness.

### v2

@@ 3.2. Empirical Findings
DeepSeek-Llama-8B 3,640 3,063 (-15.9%) +2.44% 2,568 2,564 (-0.2%) -1.85%
DeepSeek-Qwen-7B 3,839 3,464 (-9.8%) +0.61% 2,459 1,946 (-20.9%) -5.55%
DeepSeek-Qwen-14B 3,282 2,890 (-11.9%) +1.82% 2,078 1,952 (-6.1%) -1.06%
DeepSeek-Qwen-32B 3,136 2,610 (-16.8%) +1.83% 2,333 1,890 (-19.0%) +1.58%
QwQ-32B 2,363 2,236 (-5.4%) 0.00% 1,929 1,779 (-7.8%) +0.53%
Qwen3-8B 3,604 2,490 (-30.9%) +3.05% 3,663 2,545 (-30.5%) +3.70%
gpt-oss-20b 481 238 (-50.5%) +4.88% 470 241 (-48.7%) -2.64%
As shown in Table 3, introducing this prompt instruction reduces the average CoT length for most evaluated models. This confirms that CoT verbosity is sensitive to instruction-level guidance. However, the reduction rate vary substantially across models and datasets, indicating that prompt-based control lacks reliability and offers limited effectiveness in practice.
In particular, while moderate reductions are observed for several DeepSeek and Qwen models, the overall compression ratios are relatively small and highly model-dependent. In contrast, gpt-oss-20b exhibits a much more pronounced reduction in reasoning length (over 50% on HumanEval). One possible explanation is that gpt-oss-20b natively supports explicit control of reasoning length, whereas most other open-source models lack such mechanisms, making them less responsive to high-level conciseness instructions.
It should be noted that shortening CoT via prompting does not uniformly degrade performance. In several cases, accuracy remains stable or even improves after applying the concise-CoT prompt (e.g., Qwen3-8B and DeepSeek-Qwen-32B on HumanEval). This observation reinforces our Finding 3.2 that excessively long CoT does not necessarily lead to better reasoning. On the contrary, overly verbose reasoning may instead introduce redundant or noisy content that hinders effective problem-solving. Overall, prompt design can influence CoT length but lacks robustness and consistency across models. The fact that shorter CoT can maintain or even improve accuracy shows that reducing unnecessary reasoning is both feasible and useful.
    Finding 5:  Prompt-based conciseness control is limited and model-dependent. Conciseness prompts reduce CoT length for most models, but the gains are generally small and inconsistent. Only gpt-oss-20b shows a substantial reduction (over 50% on HumanEval), while accuracy is largely preserved.
@@ 4.1. Pre-Inference Data Generation
The first stage of SEER focuses on constructing a comprehensive corpus of CoT outputs to serve as the foundation for subsequent filtering and fine-tuning. In this step, the base language model is used to process each question in the training set, generating a complete solution that includes an explicit, step-by-step reasoning path. This step does not rely on prompt design. Therefore, we intentionally avoid using complex prompts. For example, for code generation tasks, we adopt the same prompt as in Section 3.1.4. To avoid early truncation that may prevent the model from completing its reasoning and based on the reasoning length observed in our empirical study, we adopt a moderate token budget during generation, i.e., 16k. Moreover, this setting can help collect more complete and diverse reasoning traces, thereby reducing information loss and mitigating overfitting caused by insufficient CoT data.
4.1.1. Best-of-N Sampling for Data Refinement
Since reasoning loops are a common phenomenon in CoT generation for reasoning models, we adopt a BoN sampling strategy to mitigate looping behaviors. In addition, this strategy enables us to collect more diverse reasoning traces for constructing a high-quality training set, and can be used to expand the dataset when the available training data are limited.
Specifically, for each input question, the model generates N candidate responses, each accompanied by a CoT. These candidates are then ranked and filtered according to a set of criteria: ❶ Correctness of the Final Answer: only candidates producing the correct final answer are considered. All incorrect outputs are discarded. ❷ Presence of Valid CoT: from the remaining candidates, we retain only those containing a non-empty and non-looping reasoning path. This guarantees that each example is explainable and preserves its interpretability. ❸ Conciseness of Reasoning: if multiple candidates satisfy the first two criteria, we select the one with the shortest CoT length. This prioritizes efficiency and prevents the inclusion of unnecessarily verbose reasoning.
4.2. Adaptive CoT Filtering
After BoN sampling, CoT lengths remain highly variable and often contain long-tail verbosity. To obtain concise yet correct training signals, we apply a two-stage filter.
Stage 1: Output validity. We first remove samples whose final outputs are invalid (e.g., code that cannot be parsed or fails the provided unit tests in code-generation datasets). This step filters out malformed or incomplete outputs before applying any length-based filtering.
Stage 2: Robust length filter. On the remaining samples, we filter overly long CoTs using a robust, distribution-aware threshold. This is motivated by the empirical observation that CoT lengths are often heavy-tailed: a small fraction of samples can be extremely long due to overthinking or looping, and these outliers can dominate both training cost and the learned behavior.
@@ 5.1. Datasets
To evaluate the performance and effectiveness of our method, we conduct experiments on a set of representative software engineering datasets. The experiments covers diverse task formats and reasoning requirements.
5.1.1. Training and evaluation benchmarks.
We use MathQA-Python (Austin et al., 2021), CodeXGLUE-Defect-Detection (Zhou et al., 2019), and Code-Search-WebQuery (Lu et al., 2021; Husain et al., 2019). All of these datasets provide predefined training and test splits. MathQA-Python focuses on generating executable Python programs to solve mathematically formulated problems, requiring both symbolic reasoning and correct code implementation. CodeXGLUE-Defect-Detection is a binary classification task that assesses whether a given code snippet contains functional defects or security vulnerabilities, emphasizing program understanding and static reasoning. Code-Search-WebQuery evaluates semantic code retrieval by determining whether a code snippet satisfies a natural language query, challenging the model to align program semantics with textual descriptions. These three datasets are used to train our method and all baseline models under identical settings, with evaluation conducted on their respective test sets.
5.1.2. Evaluate-only benchmarks.
In addition, we evaluate our method on HumanEval and MBPP-Sanitized, which are used exclusively for evaluation. These datasets are not involved in training, and serve to assess the generalization capability of our method on unseen code generation tasks.
Overall, this benchmark suite enables a systematic evaluation of both in-domain performance and cross-task generalization across software engineering scenarios.
@@ 5.2. Implementation Details
All experiments are conducted on a server equipped with four NVIDIA RTX A800 GPUs (80 GB each) and 128 Intel Xeon CPUs, running Ubuntu 22.04 LTS. We fine-tune DeepSeek-R1-Distill-Qwen-7B model (DeepSeek-AI, 2025) under a full-parameter supervised fine-tuning (SFT) setting, using the LLaMAFactory framework (Zheng et al., 2024). Training is performed for three epochs with an effective batch size of eight, gradient checkpoint enabled, and the AdamW optimizer is used. The learning rate is fixed at $1\times 10^{-5}$ with a cosine decay schedule and 10% warm-up, and gradient clipping (norm 1.0) is applied for stability. During fine-tuning, inputs are tokenized with a maximum sequence length of 16,384 tokens. Although we adopt full-parameter SFT by default, the SEER framework is also compatible with parameter-efficient fine-tuning methods such as LoRA and QLoRA, which we discuss in the Discussion section.
@@ 5.3. Baselines
  • Short CoT: Employ prompting techniques to elicit a more concise CoT from the model, using the prompt design described in Section 3.
@@
  • Naive BoN (Munkhbat et al., 2025): Use only the Best-of-N method to process the pre-inference data and then train the model.
@@ 5.4. Evaluation Metrics
Pass@1: We report pass@1 as the primary metric to evaluate correctness, i.e., whether the model produces a correct output on the first attempt. For programming tasks (e.g., MathQA-Python), a prediction is considered correct if the generated code passes all test cases. For binary classification tasks (e.g., CodeXGLUE-Defect-Detection), a prediction is considered correct if the predicted label matches the ground-truth label, and the resulting accuracy is equivalent to pass@1 under single-attempt evaluation.
Compression Rate: We define the Compression Rate $\rho$ to quantify the conciseness of the generated CoT. Specifically, $\rho$ measures the relative reduction in the number of tokens compared to a reference CoT. It is computed as $\rho=1-(\lambda^{\prime}/\lambda)$. In this formula, $\lambda$ denotes the token count of the reference (original) CoT, and $\lambda^{\prime}$ represents the token count of the CoT generated by the evaluated model or baseline. A higher value of $\rho$ indicates a greater degree of compression, corresponding to a more concise reasoning path.